OpenStudy (anonymous):

I'm using De Moivre's theorem to get all the solutions for z^4 =i but I'm unsure how to get the angle theta. Do I use r or the polar form?

OpenStudy (anonymous):

is that supposed to be z^4 = 1? or i?

OpenStudy (anonymous):

SEE \[\phi = \tan^{-1} y/x\]

OpenStudy (anonymous):

Z^4 = i If z = x +iy so x is zero and y is 1

OpenStudy (anonymous):

\[e ^{i \theta} \]You can work it now ..

OpenStudy (anonymous):

\[y/x \rightarrow \infty\]\[\theta = 90\]

OpenStudy (anonymous):

hmmm....i really think it should be: \[x^{4} = 1\] The fourth roots of unity. If you use wolfram with the equation: \[x^{4} = i\] you get some pretty bizarre answers. and yeah, \[e^{i\theta}\] is used with the nth roots of unity, where: \[\theta = \frac{2\pi}{n}\] so in this case, since n = 4 \[\theta = \frac{2\pi}{4} = \frac{\pi}{2}\] That would be the angle hes looking for.

OpenStudy (anonymous):

same answers lol, now we know its right.

OpenStudy (anonymous):

Thanks